The risk-free rate of return is 6%, the required rate of return on the market is 12%, and High-Flyer stock has a beta coefficient of 1.0. If the dividend per share expected during the coming year, D1, is $2.00 and g = 4%, at what price should a share sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Share price
Answer #1
1] Required return per CAPM = risk free rate+beta*(expected return on the market-risk free rate)
= 6%+1*(12%-6%) = 12.00%
2] Price per the constant dividend growth model = D1/(r-g) = 2/(0.12-0.04) = $          25.00